OpenStudy (moazzam07):

Sam bought 86 pounds of flour for his bakery. Every day he used the same amount of flour to make bakery items. After 10 days, Sam was left with 25.2 pounds of flour. On average, how many pounds of flour did Sam use each day? 3.01 pounds 5.65 pounds 6.08 pounds 11.12 pounds

OpenStudy (moazzam07):

I need help on this i dont get it

OpenStudy (moazzam07):

i men i get but i dont know how to do it

OpenStudy (kash_thesmartguy):

86-10x=25.2 Solve for x.

OpenStudy (moazzam07):

how

OpenStudy (taramgrant0543664):

All you have to do is take the initial weight minus the weight after the 10 days and divide by the number of days this will show approximately how much flour was used each day

OpenStudy (kash_thesmartguy):

First, subtract 86 from both sides. \[(86-86)+10x=25.2-86.\]

OpenStudy (moazzam07):

x= 6.08?

OpenStudy (moazzam07):

or 68?

OpenStudy (taramgrant0543664):

6.08 is the answer!

OpenStudy (kash_thesmartguy):

It would be 6.08

OpenStudy (moazzam07):

i knew it

OpenStudy (moazzam07):

thanks @Kash_TheSmartGuy and @taramgrant0543664

OpenStudy (kash_thesmartguy):

But @taramgrant0543664 , he needs an equation to further explaining

OpenStudy (study_together):

It is 6.08 pounds each day

OpenStudy (kash_thesmartguy):

First, subtract 86 from both sides. \[(86-86)+10x=25.2-86.\] do this

OpenStudy (moazzam07):

ok you want me to 25.2 - 86 = -60.8

OpenStudy (study_together):

first you subtract 25.2 from 86. you get 60.8. you do 60.8/10

OpenStudy (moazzam07):

yes i solved it

OpenStudy (moazzam07):

I get it now

OpenStudy (kash_thesmartguy):

Good! Now do this\[\frac{ 10x }{ 10 }=\frac{ -60.8 }{ 10 }\]

OpenStudy (kash_thesmartguy):

Good job! @moazzam07

OpenStudy (taramgrant0543664):

@Kash_TheSmartGuy your equation is wrong you should have 25.2+10x=86

OpenStudy (moazzam07):

60.8 / 10 = 6.08

OpenStudy (study_together):

and you have your answer

OpenStudy (kash_thesmartguy):

No @taramgrant0543664 as the equation will result in 25.2 instead of start with it

OpenStudy (taramgrant0543664):

@Kash_TheSmartGuy -60.8/10 is a negative number not the positive number you are looking for

OpenStudy (kash_thesmartguy):

Sorry, there's a problem. In my equation it should be divided by -10 instead of 10.

OpenStudy (taramgrant0543664):

It could also be 86-10x=25.2 but you had a + after the 86

OpenStudy (kash_thesmartguy):

So, you get a positive answer. Final Answer: 6.08